(1) It took her 9 more months but Marina has managed to save the full $725...

(1)

It took her 9 more months but Marina has managed to save the full $725 plus more to cover fees to pay off the pay-day loan company. However, she forgot to account for the interest that had been compounding over time. Consider it is now 275 days later, the remaining loan was $725 and the APR is 47% compounded daily.  

What is the total amount that Marina must now pay in order to pay off her the loan, accounting for interest?

What is the total amount of interest paid (not including fees)?

(2)

Tony is saving up to go traveling after graduation. How much will he have if he can put $100 at the beginning of every month, for the full four years he is in school, into his high-interest savings account which earns 1.95% interest?

How much interest will be earned?

(3)

Anna also wants to travel after graduation but she isn’t working while in school so she won’t be able to put money aside regularly. However, her grandmother gave her the generous gift of $2500 for her birthday and suggested she invest it in a GIC because it’s low risk. She shopped around and found a 4 year GIC investment that earns 2.8% interest annually.

How much will Anna have at the end of her four year investment?

How much interest will be earned?

How much interest will be earned due to compounding*?

*calculate simple interest using the simple interest formula, then compare to the total interest earned to discover the impact of compounding the interest
Answer #1

(1) Loan Outstanding = $ 725 exlcuding fees, Interest Rate = 47% APR compounded daily

Daily Interest Rate = (47/365) ~ 0.1288%

Total Amount Outstanding after 275 days = 725 x (1.001288)^(275) = $ 1032.825

Amount Required to payoff the loan is ~ $ 1032.825

Interest Accumulated = Total Required Payoff - Loan Outstanding = 1032.825 - 725 = $ 307.825

NOTE: Please raise separateq queries for solutions to the remaining unrelated questions as one query is restricted to the solution of only one complete question (with a maximum of four sub-parts)